As at the 1/1/2004 stock of finished goods at market value GHS 4200 (in units 300

transferred at cost plus 20%) during the year unit started were 4000 but completed

units were 3800 and were transferred to warehouse at cost plus 20%. Total direct

cost and total indirect cost on all the 4000 units at production centers were GHS85,

000 and GHS45,000 respectively. Opening stock of work-in-progress was 400

units valued at the factory cost of GHS6, 000, and the value of the work-in-progress

at the close of the year was GHS2000.unit sold during the year was 3300

respectively. You are required to determine the following:

A. Manufacturing profit

i. Total Manufacturing profit realized during the period

ii. Total Manufacturing profit created during the period

iii. Total Manufacturing profit created during the period and realized




1
Expert's answer
2020-07-20T18:07:11-0400


Let the selling price be 100.

Then

i."(3300\times 100+(4200)-(4200*0.2+4200))-(85000+45000)=(330 000+4200-5040)-(130 000)=199 160"

ii.

"(3300\times 100+6000-2000-(85000+45000)=334 000-(130 000)=204 000"

iii.

"(3300\times 100+(4200)-(4200*0.2+4200))+6000-2000-(85000+45000)=(330 000+4200-5040+6000-2000)-(130 000)=203 160"